# Reasoning with Confidence: Efficient Verification of LLM Reasoning Steps via Uncertainty Heads

### v1

@@ Abstract
Solving complex tasks usually requires LLMs to generate long multi-step reasoning chains. Previous work has shown that verifying the correctness of individual reasoning steps can further improve the performance and efficiency of LLMs on such tasks and enhance solution interpretability. However, existing verification approaches, such as Process Reward Models (PRMs), are either computationally expensive, limited to specific domains, or require large-scale human or model-generated annotations. Thus, we propose a lightweight alternative for step-level reasoning verification based on data-driven uncertainty estimation. We train transformer-based uncertainty quantification heads (UHeads) that use the internal states of the frozen LLM to estimate the uncertainty of its reasoning steps during generation. The approach is fully automatic: target labels are generated either by another larger LLM (e.g., DeepSeek R1) or in a self-supervised manner by the original model itself. UHeads are both effective and lightweight, containing less than 10M parameters. Across multiple domains, including mathematics, planning, and general knowledge question answering, they match or even exceed the performance of PRMs that are up to 810× larger. Our findings suggest that the internal states of LLMs encode their uncertainty and can serve as reliable signals for reasoning verification, offering a promising direction towards scalable and generalizable introspective LLMs. Equal contribution. Correspond to {jingni,efadeeva,msachan}@ethz.ch, tianyi_wu@u.nus.edu, artem.shelmanov@mbzuai.ac.ae
@@ 1 Introduction
Chain-of-thought (CoT) prompting has proven highly effective in eliciting the reasoning capabilities of large language models (LLMs) for solving complex tasks (wei2022chain). Recent post-training approaches further enhance this ability through reinforcement learning, rewarding models for generating responses that conform to the CoT pattern and yielding correct final answers (deepseekai2025deepseekr1incentivizingreasoningcapability; yang2025qwen3; abdin2025phi). However, the emphasis on verifying only final answers raises concerns about the reliability of intermediate reasoning steps (lightman2023letsverifystepstep; zheng2025processbenchidentifyingprocesserrors; Barez2025_CoT_not_explainability). Flawed steps can propagate and systematically distort conclusions (fu2025deepthinkconfidence), while in some cases LLMs may still produce the correct answer despite erroneous intermediate steps (arcuschin2025chainofthoughtreasoningwildfaithful). Such issues undermine trust in CoT-based methods, particularly in high-stakes domains like medicine (AmannEtAl2020_Explainability_AI_Healthcare) and law (fan2025lexambenchmarkinglegalreasoning).
Figure 1: Process reward models (top) in comparison to uncertainty quantification heads (bottom).
@@
The contemporary approach to supervise models for step-level correctness is through process reward models (PRMs: lightman2023letsverifystepstep; zhang2025lessonsdevelopingprocessreward). However, PRM-based supervision faces key drawbacks. First, it requires expensive monte-carlo rollouts for training data annotation and deploying an additional LLM supervisor, adding substantial computational overhead. Second, this approach fails in tasks like mathematical proofs (azerbayev2023proofnetautoformalizingformallyproving) and planning (zheng2024naturalplanbenchmarkingllms), where step-level correctness cannot be inferred from the final result alone.
Another promising line of work for assessing the correctness of reasoning steps leverages uncertainty quantification (UQ) methods (gal2016uncertainty; malinin2020uncertainty). Unlike PRMs, which rely on external verification, UQ assumes that a model’s outputs and internal states provide information about the reliability of its generations. To date, only relatively simple unsupervised UQ methods have been investigated for LLM reasoning (fu2025deepthinkconfidence; yan2025mur). These approaches are primarily based on consistency-based heuristics such as majority voting and rely on simplistic uncertainty estimates without leveraging any model internals. While computationally lightweight, they consistently fail to capture subtle reasoning flaws and lag significantly behind PRMs in performance.
In this work, we ask: Can the cost-efficiency of UQ methods be combined with the performance of PRMs? Recent research in UQ shows that data-driven approaches substantially outperform unsupervised ones in detecting factual hallucinations (azaria-mitchell-2023-internal; chuang-etal-2024-lookback; shelmanov2025headpredictheadquestion). Building on this insight, we introduce a lightweight data-driven uncertainty-based verifier for step-level reasoning. Specifically, we adapt the UHead framework (shelmanov2025headpredictheadquestion) to train uncertainty quantification heads on top of frozen LLMs for reasoning step verification. Unlike PRMs, which rely on expensive rollouts and training of a large supervisor LLM, UHeads exploit internal LLM states (attention and token probabilities) as features, making them far more efficient (see Figure 1): UHeads trained in our work require less than 10M parameters. Training labels are generated automatically, either by using a powerful external LLM (e.g., DeepSeek R1) as a judge or by relying on the original LLM itself. This enables scalable supervised or self-supervised training without the need for verifiable final answers, human annotations, or costly Monte Carlo rollouts.
@@
Extensive in-domain (ID) and out-of-domain (OOD) evaluation highlights UHeads’ advantages in three key aspects: (1) Generalizable Performance: UHeads achieve competitive or superior performance compared to PRM baselines in step-level verification and test-time scaling, including best-of-$N$ selection and step-level correctness supervision, especially when dealing with OOD reasoning tasks (e.g., reasoning-intensive QA and planning); (2) Computational Efficiency: UHeads substantially outperform PRMs that are 150 larger, while remaining competitive with PRMs up to 800 larger; and (3) Training Data Efficiency: unlike PRMs, which often rely on proprietary datasets (skyworkopeno12024), costly human annotations (lightman2023letsverifystepstep), or consensus-filtering pipelines involving both Monte Carlo rollouts and LLM judgments (zhang2025lessonsdevelopingprocessreward), UHeads can be trained on data annotated automatically in a self-supervised manner.
Our contributions are the following:
  • We introduce a lightweight approach to verifying step-level correctness in LLM-generated reasoning by extending the UHead framework. Our 9.8M-parameter UHead achieves competitive or superior performance compared to 7B-scale PRMs across diverse reasoning tasks and settings, while being significantly more efficient at inference time in terms of memory and compute.
  • We develop an efficient and scalable data annotation pipeline that eliminates the need for human-labeled data, verifiable final answers, or costly Monte Carlo rollouts. We further show that high performance in reasoning verification can be achieved using annotations derived either from external verification models (e.g., DeepSeek R1) or from a fully self-supervised process.
  • We show that our approach achieves strong generalization across domains: a single UHead trained only on mathematical tasks generalizes effectively to planning and general knowledge QA. Unlike PRMs, which often overfit to domain-specific reasoning patterns, our lightweight UHead verifier leverages internal LLM uncertainty signals, enabling robust reasoning verification without requiring domain adaptation.
@@ 2.1 Test-Time Scaling
State-of-the-art LLMs solve complex tasks by generating CoT reasoning traces (guo2025deepseek). Given an input question $\mathbf{x}$, models produce a sequence of reasoning steps $\mathbf{r}=\{r_{1},r_{2},\dots,r_{T}\}$ followed by a final answer $\mathbf{y}=g(\mathbf{r},\mathbf{x})$. Formally, an LLM parameterized by $\theta$ models the conditional distribution:
[…]
where $r_{t}$ denotes the $t$-th reasoning step and $\mathbf{r}_{<t}$ are the preceding steps. Although CoT reasoning improves final task performance, locally incorrect steps $r_{t}$ can propagate and cause wrong answers.
@@
Test-time scaling (TTS) algorithms (yao2023tree; snell2024scaling) aim to further improve the performance of LLMs without retraining the entire model, by allocating additional computation or using more sophisticated inference strategies. By leveraging multiple candidate solutions and verification signals, TTS facilitates selecting a correct answer. Common approaches include online and offline best-of-$N$ (BoN) sampling. Other techniques employ sophisticated search and sampling techniques based on beam search and backtracking, such as the tree of thoughts (yao2023tree).
Online best-of-$N$ sampling assumes that in each step, we generate $N$ continuations $\{r_{t}^{(1)},r_{t}^{(2)},\dots,r_{t}^{(N)}\}$, and each of them is evaluated online using a quality function $Q_{\text{online}}\!\left(\mathbf{r}^{(j)}_{1:t}\right)$. In the greedy version of online best-of-$N$, only the most promising option is expanded further. After reaching the final step $T$, the answer is obtained from the greedily derived chain:
$r^{*}_{t}=\arg\max_{j\in\{1,\dots,N\}}Q_{\text{online}}\!\left(\mathbf{r}^{(j)}_{1:t}\right),\;\mathbf{r}^{*}=\{r_{1}^{*},r_{2}^{*},...,r_{T}^{*}\},\;\mathbf{y}=g(\mathbf{x},\mathbf{r}^{*}).$
@@
Offline best-of-$N$ sampling assumes that we sample not one, but $N$ reasoning chains $\{\mathbf{r}^{(1)},\mathbf{r}^{(2)},\dots,\mathbf{r}^{(N)}\}$. Each chain $\mathbf{r}^{(j)}$ is evaluated using a scoring function $Q_{\text{offline}}\!\left(\mathbf{r}^{(j)}\right)$ (see Section 2.2), and the final answer is derived from the best-scoring chain:
$\mathbf{r}^{*}=\arg\max_{j\in\{1,\dots,N\}}Q_{\text{offline}}\!\left(\mathbf{r}^{(j)}\right),\quad\mathbf{y}=g(\mathbf{x},\mathbf{r}^{*}).$
2.2 Process Reward Models
@@
PRMs are a prominent approach for quality assessment in both online and offline test-time scaling (lightman2023letsverifystepstep; luo2024improvemathematicalreasoninglanguage; wang-etal-2024-math). PRMs are critics designed to estimate the quality of each partial reasoning state $\mathbf{r}^{(j)}_{1:t}$ by assigning a reward $R_{\text{PRM}}\!\left(\mathbf{r}^{(j)}_{1:t}\right)$ that reflects the likelihood of the chain eventually leading to a correct solution. PRMs are typically implemented by a separate LLM trained to evaluate the plausibility and correctness of intermediate reasoning steps. The annotation for training PRMs comes from various sources, including crowdsourcing, self-consistency checking, and synthetic data generation pipelines.
@@
In online BoN, we simply use the reward from PRM as a step-wise quality function. In offline BoN, the score for the complete chain $\mathbf{r}^{(j)}$ can be obtained via a temporal aggregation of process rewards, e.g., minimum step score (zhang2025lessonsdevelopingprocessreward):
$Q_{\text{online}}\!\left(\mathbf{r}^{(j)}_{1:t}\right)=R_{\text{PRM}}\left(\mathbf{r}^{(j)}_{1:t}\right),\quad Q_{\text{offline}}\!\left(\mathbf{r}^{(j)}\right)\;=\;\min_{1\leq t\leq T^{(j)}}R_{\text{PRM}}\!\left(\mathbf{r}^{(j)}_{1:t}\right)$.
@@
Although PRMs have proven to be highly effective even in boosting the performance of larger LLMs, they come with several limitations: they are relatively large models, typically containing 1.5B–8B parameters, which leads to additional computational and memory overhead during inference. Moreover, PRMs are often domain-specific, being trained for tasks such as mathematical reasoning, and exhibit limited generalization to unseen domains.
2.3 Uncertainty Quantification
UQ is a conceptual and methodological framework that bridges Bayesian and information-theoretical approaches to modeling (maddox2019simple) with various practical tasks such as selective classification (geifman2017selective), out-of-distribution detection (mukhoti2023deep), and adversarial attack detection (lee2018simple). The core assumption behind UQ is that ML models can produce not only predictions but also signals indicative of their reliability. In practice, it is common to use various heuristics and approximations, which can be of various natures. However, unlike external verification methods, UQ primarily leverages internal model capabilities to provide these estimates, such as logits, multiple sampled predictions, hidden states, and attention weights (see Figure 1).
UQ for LLMs poses unique challenges beyond standard text classification, making it difficult to establish an effective universal unsupervised solution (zhang-etal-2023-enhancing-uncertainty; KuhnGF23; duan-etal-2024-shifting; fadeeva2024fact). These limitations have motivated the exploration of data-driven approaches, which have consistently demonstrated superior empirical performance (azaria-mitchell-2023-internal; he-etal-2024-llm; chuang-etal-2024-lookback).
3 Verification of Reasoning Steps with Uncertainty Quantification Heads
Method. UQ offers an efficient alternative to PRMs for assessing the quality of reasoning steps during online test-time scaling. While PRMs leverage generated sequences, UQ methods leverage internal capabilities of LLMs. Instead of training large reward models, it is possible to use lightweight UQ methods or train lightweight data-driven uncertainty estimators:
[…]
We adopt a state-of-the-art data-driven method for uncertainty quantification (UQ) – uncertainty quantification heads (UHeads) (shelmanov2025headpredictheadquestion). UHeads are lightweight supervised auxiliary modules for LLMs, designed to estimate uncertainty and detect hallucinations. Implemented as compact transformer-based classifiers, they leverage the internal states of the base LLM without requiring fine-tuning or modification of the model or its outputs. Unlike PRMs, which are domain-specific (e.g., math proofs or planning) and incur significant additional inference costs, UHeads are plug-and-play modules that generalize across tasks, domains, and languages, making them practical for reasoning step assessment.
While PRMs rely solely on generated tokens, UHead leverages richer per-token features extracted from the base LLM’s internal states, specifically, attention weights to the 1-3 preceding tokens and the logits of the top-K candidate generations. Its architecture is transformer-based: each per-token feature vector is first linearly projected and then processed through a stack of $L$ Transformer blocks to capture intra-step dependencies. The resulting token representations are mean-pooled across the reasoning step to obtain a step-level vector, which is passed through a two-layer classification head with dropout and a GeLU activation to produce the final class logits.
The architecture of UHead consists of three main components:
@@ 3 Verification of Reasoning Steps with Uncertainty Quantification Heads
    Feature projection layer: a linear transformation that maps the input features:
@@
    Transformer layers: a stack of $L$ Transformer blocks:
@@
    Aggregation of token-level features into a step-level vector:
[…]
@@
    Two-layer classification head: applies two linear layers with non-linearity and dropout:
    $\mathbf{z}=\mathrm{Linear}_{2}\!\Big(\sigma\big(\mathrm{Dropout}(\mathrm{Linear}_{1}(\mathbf{h}_{L}))\big)\Big),$
    where $\sigma(\cdot)$ is an activation function (GeLU) and $\mathbf{z}$ are the class logits.
Originally developed to score atomic factual claims (fadeeva2024fact), here we extend UHeads to score reasoning steps in a PRM-style setup.
@@
Constructing training dataset for UHead. To construct the training data for UHead, we use 10.8K problems (prompts) from the PRM800K training dataset. This dataset is an established resource for training and evaluating LLM reasoning capabilities. As most PRMs are already trained on LLM generations derived from it, it enables a fair comparison (wang2024mathshepherdverifyreinforcellms; zhang2025lessonsdevelopingprocessreward).
We prompted reasoning LLMs to generate each step of the CoT solution on a separate, self-contained line, enabling accurate correctness grading. The full prompt is provided in Appendix C.1. Qwen3-8B reliably follows this prompt and produces high-quality CoT steps. For each of the 10.8K math problems, we generate 3 reasoning chains, resulting in ~32K data samples. For all reasoning chain generations, we use nucleus sampling with a maximum of $256$ new tokens, top-$k=50$, top-$p=0.95$, and a temperature of $1.0$.
@@
Then, we annotated step-level correctness of each CoT step using an LLM as a judge. The judge LLM was provided with the question, the target LLM’s CoT steps and final answer, and the ground-truth answer; and then prompted to assess the correctness of each step. The full annotation prompts are given in Appendix C.2. We consider two judging approaches: (1) an external verifier, where a larger LLM evaluates the steps, and (2) a self-supervised setting, where the same reasoning LLM annotates its own generations. Following zheng2025processbenchidentifyingprocesserrors, who showed that general reasoning models perform well at step-level correctness annotation, we adopt DeepSeek-R1 as the external verifier. Other details to reproduce the training data can be found in § D.1.
4 Experimental Setup
4.1 LLMs for Reasoning, Uhead Training Details, Evaluation Datasets
LLMs for reasoning. We conduct experiments with two state-of-the-art LLMs, Qwen3-8B (qwen3) and Phi-4 (abdin2025phi). For these models, we prepared training datasets for reasoning-step verification and trained UHeads on them.
UHead training settings. We follow shelmanov2025headpredictheadquestion’s recipe to train UHeads with a frozen underlying LLM. For all experiments with different training data and underlying LLMs, we use the same hyperparameter settings, which are detailed in § D.2.
Evaluation datasets span three domains: mathematical reasoning (in-domain), planning (OOD), and general knowledge QA (OOD). We select datasets that demand non-trivial reasoning and include reasoning steps that are unambiguously verifiable, enabling reliable evaluation by both LLMs and human annotators. The details for test datasets are presented in Appendix E.
4.2 Baselines
Small PRMs (1.5B) include 2 PRMs fine-tuned from Qwen2.5-Math-1.5B: Skywork-PRM-1.5B (skyworkopeno12024) and H4-Qwen2.5-PRM-1.5B-0.2 (huggingfaceh4_qwen2.5_math1.5b_instruct_prm0.2).
Large PRMs (7–8B) include (1) Math-Shepherd-PRM-7B (wang-etal-2024-math), which determines the process labels for each step by estimating the empirical probability of reaching the correct final answer (MC estimation); (2) RLHFlow-PRM-8B-DeepSeek/Mistral models (xiong2024implementation) trained with MC-estimated labels from DeepSeek/Mistral rollouts; (3) Universal-PRM-7B (auroraprm2025) trained using ensemble prompting and reverse verification; (4) Qwen2.5-Math-7B-PRM800k trained on the PRM800k dataset (lightman2023letsverifystepstep); and (5) Qwen2.5-Math-PRM-7B (zhang2025lessonsdevelopingprocessreward), which combines MC estimation with LLM-as-a-Judge consensus and currently achieves the best result on ProcessBench compared to PRMs of similar scale and computation (zhao2025genprm).
UQ methods evaluated in our experiments fall into two categories, differing in computational cost: (1) lightweight scores that use only single generation: Maximum Sequence Probability, Mean Token Entropy, Perplexity (fadeeva-etal-2023-lm), P(True) (kadavath2022languagemodelsmostlyknow), CCP (fadeeva2024fact), and Self-Certainty (kang2025scalable), which show significant advantages in reasoning chain selection (fu2025deepthinkconfidence); and (2) sampling-based scores: Semantic Entropy, Lexical Similarity (kuhn2023semantic), and Degree Matrix (lin2024generating). To compute the latter, we draw $M=10$ alternative steps per step position, so they are much less computationally efficient. All methods are implemented using the LM-Polygraph framework (fadeeva-etal-2023-lm; vashurin2024benchmarking).
4.3 Evaluation Settings
Dataset # Steps Acc.
PRM800K 17,067 95.29
ProofNet 193 87.05
Trips 102 93.06
Meetings 101 95.70
Calendar 102 91.09
StrQA 313 95.85
SciQA 245 99.28
We conduct experiments to compare and evaluate PRMs and UQ methods in three scenarios: (1) step-level correctness prediction; (2) offline best-of-$N$ selection; and (3) online best-of-$N$ selection.
(1) Step-level correctness prediction. For each question in the test set, we generate a CoT trace and evaluate step-level correctness using DeepSeek-R1. To ensure accurate judgements, the model is provided with full metadata, such as the ground-truth answer, reasoning steps, and supporting evidence (see prompt in § C.2). The resulting correctness labels serve as the reference for evaluating the step-level quality scores from our baselines. To validate the reliability of our evaluation pipeline, we evaluate it against (1) the human annotations from a random subset of PRM800k and (2) a manually annotated set of ~1000 steps spanning QA, planning, and ProofNet tasks (see Appendix G for annotation details). DeepSeek-R1 achieves 95% Acc. on PRM800k and ~90% on other datasets (see Table 1).
(2) Offline best-of-$N$. In the Best-of-$N$ (BoN) setting, we generate $N$ reasoning chains per problem (N=10N{=}10 for the math and QA datasets, N=5N{=}5 for the planning datasets, temperature=1.0). We compute the quality score $Q$ for each step, and the chain with the highest score is selected for deriving the final answer. Correctness of the solution is measured using accuracy. For GSM8K, we use exact match against the gold-standard final answer. For other datasets, where final answers may be open-ended or structurally complex, we use DeepSeek-R1 to assign binary correctness labels (1 for correct, 0 for incorrect) based on both the problem statement and the reference solution. The grading prompt is provided in Appendix C.2.
(3) Online best-of-$N$ selects the best candidate at each step during the generation process. At each step, the system produces $N$ candidate steps, scores them, and expands the best option. This process repeats until an end-of-sequence token is reached, after which the chain is evaluated using the same metrics as in the offline BoN. This setting evaluates how well PRMs and UQ methods can guide the LLM reasoning trajectory. For computational efficiency, we pick N=5N{=}5 and use a generation temperature of 1.5 to ensure the diversity of candidate steps.
[…]

### v3

# Efficient Test-Time Scaling of Multi-Step Reasoning by Probing Internal States of Large Language Models

@@ Abstract
[…]
Efficient Test-Time Scaling of Multi-Step Reasoning by
Probing Internal States of Large Language Models
[…]
@@ 1 Introduction
Chain-of-thought (CoT) prompting has proven highly effective in eliciting the reasoning capabilities of Large Language Models (LLMs) to solve complex tasks (Wei et al., 2022). Recent post-training approaches further enhance this capability through reinforcement learning, where models are rewarded for producing responses that demonstrate chain-of-thought reasoning prior to emitting the final answer, leading to the range of latest Language Reasoning Models (LRMs) (DeepSeek-AI, 2025; Yang et al., 2025; Abdin et al., 2025).
Even the best LRMs generate incorrect reasoning trajectories and ultimately erroneous outputs. A single flawed reasoning step can derail the entire solution. Empirical evidence suggests, however, that allowing the model to attempt the task multiple times or exploring multiple candidate continuations at intermediate steps can increase the likelihood of reaching a correct answer – a paradigm known as test-time scaling (TTS) (Yao et al., 2023; Snell et al., 2024). Many TTS strategies rely on a scorer that evaluates the quality of individual reasoning steps and provides guidance to steer the reasoning process towards better solutions.
Figure 1: Left: Given a problem set, target LLM generates CoTs and an annotator LLM labels step-level correctness. Middle: Training ReProbe with target LLM providing internal states while frozen. Right: at inference, ReProbe and PRM monitors internal states and output text correspondingly. ReProbe has significantly fewer parameters.
@@
The contemporary approach to reasoning step verification is through process reward models (PRMs: Lightman et al., 2024; Zhang et al., 2025c). Originally designed to guide the RL post-training, PRMs were found to be particularly useful for TTS as well. However, PRM-based verification faces key drawbacks. First, it requires expensive Monte-Carlo rollouts for training data annotation and a significant amount of computation for fine-tuning. Constructing training data is also complicated in tasks where step-level correctness cannot be inferred from the final result, such as mathematical proofs (Azerbayev et al., 2023) and planning (Zheng et al., 2024). Second, deploying PRMs incurs substantial inference-time overhead, as they require running an additional LLM that significantly increases GPU memory usage and computational cost. Third, PRMs are typically fine-tuned on narrow domains, such as mathematics, and therefore exhibit limited generalization.
In another interesting line of work, the correctness of reasoning steps is assessed through uncertainty quantification (UQ) methods (Gal and others, 2016; Malinin and Gales, 2021). Unlike PRMs, which represent external knowledge, UQ assumes that a model’s outputs and internal states provide information about the reliability of its generations (Fu et al., 2025; Yan et al., 2025). While computationally lightweight, unsupervised UQ methods typically exhibit limited effectiveness in related tasks such as hallucination detection (Vazhentsev et al., 2025; Shelmanov et al., 2025).
In this work, we ask: can the cost-efficiency of UQ methods be combined with the performance of PRMs? Prior work in UQ shows that LLM internal states encode rich credibility signals that can be efficiently extracted using learned classifiers for hallucination detection (Azaria and Mitchell, 2023; Su et al., 2024; Shelmanov et al., 2025). Building on this insight, we introduce a lightweight training-based Reasoning Probe (ReProbe) that uses LLM internal states to determine the credibility of generated reasoning steps. Unlike PRMs, which rely on external knowledge to evaluate generated reasoning traces, ReProbe operates directly on LLM internal signals produced during generation – hidden states, attention weights, and logits – making step-level reasoning verification substantially more efficient (see Fig 1).
@@
Extensive in-domain (ID) and out-of-domain (OOD) evaluations highlight the advantages of our approach in three key aspects: (1) Strong generalizable performance: ReProbes achieve competitive or superior performance compared to PRMs in step-level verification and test-time scaling, especially when dealing with OOD reasoning tasks; (2) Computational efficiency: ReProbes outperform PRMs that are up to $150\times$ larger and remain competitive with PRMs up to $800\times$ larger. While typical PRMs contain 1.5–8B parameters, the probes trained in our work require fewer than 10M parameters; (3) Training data efficiency: unlike PRMs, which often rely on proprietary datasets (He et al., 2024), costly human annotations (Lightman et al., 2024), or consensus-filtering involving both Monte Carlo rollouts and LLM judgments (Zhang et al., 2025c), ReProbes can be trained on data annotated automatically in a self-supervised and cost-efficient manner.
@@ 2.1 Test-Time Scaling
Figure 2: Illustration of the two prevalent test-time scaling methods: Best-of-N and Beam Search.
@@
LLMs solve complex tasks by generating CoT reasoning traces (Guo et al., 2025). Given an input question $\mathbf{x}$, models produce a sequence of reasoning steps $\mathbf{r}=\{r_{1},r_{2},\dots,r_{T}\}$ followed by a final answer $\mathbf{y}=g(\mathbf{r},\mathbf{x})$, and we assume that we can effectively extract individual steps and the final answer. Although linear CoT reasoning improves final task performance, locally incorrect steps $r_{t}$ can propagate and cause incorrect answers.
@@
Test-time scaling (Yao et al., 2023; Snell et al., 2024) aims to further improve the performance of LLMs without retraining the entire model by allocating additional computation or using more sophisticated inference strategies. By leveraging multiple candidate solutions and verification signals, TTS facilitates the selection of a correct answer. Common approaches include best-of-$N$ (BoN) sampling and beam search (Xie et al., 2023; Yao et al., 2023; Snell et al., 2025) (Fig 2).
@@
Best-of-NN (BoN) assumes that we sample not one, but $N$ reasoning trajectories $\{\mathbf{r}^{(1)},\mathbf{r}^{(2)},\dots,\mathbf{r}^{(N)}\}$. Each trajectory $\mathbf{r}^{(j)}$ is evaluated using a scoring function $Q_{\text{BoN}}\!\left(\mathbf{r}^{(j)}\right)$, and the final answer is derived from the best-scoring trajectory:
$\mathbf{r}^{*}=\arg\max_{1\leq j\leq N}Q_{\text{BoN}}\!\left(\mathbf{r}^{(j)}\right),\quad\mathbf{y}=g(\mathbf{x},\mathbf{r}^{*}).$
Beam Search (BS) (Xie et al., 2023; Snell et al., 2025), also known as the tree-of-thought with the breadth first search (Yao et al., 2023), assumes that we keep $B\geq 1$ best reasoning trajectories so far; and for each of them, at a new step, we sample $N$ variants of continuations: $\{r_{t}^{(i,1)},r_{t}^{(i,2)},\dots,r_{t}^{(i,N)}\},1\leq i\leq B$. Each of the continuations is assessed using a quality function $Q_{\text{bs}}\!\left(\mathbf{r}^{(i,k)}_{1:t}\right)$, and reasoning trajectories are expanded with the best continuations $\mathbf{r}^{(i)}_{1:t}=\mathbf{r}^{(i)}_{1:t-1}\circ r_{t}^{(i,*)}$. After reaching the final step $T$, the answer is obtained from the best reasoning trajectory.
$S^{\prime}_{t}=\left\{\,\mathbf{r}^{(i)}_{1:t-1}\circ r_{t}^{(i,k)}\;\middle|\;\mathbf{r}^{(i)}_{1:t-1}\in S^{*}_{t-1},\;1\leq k\leq N\right\},$
$S^{*}_{t}=\arg\max_{S\subseteq S^{\prime}_{t},\;|S|=B}\sum_{\mathbf{r}^{(i,k)}_{1:t}\in S}Q_{\text{bs}}\!\left(\mathbf{r}_{1:t}^{(i,k)}\right),$
$\mathbf{r}^{*}=\arg\max_{\mathbf{r}^{(i)}_{1:T}\in S^{*}_{T}}Q_{\text{bs}}\left(r^{(i)}_{1:T}\right),\;\mathbf{y}=g(\mathbf{x},\mathbf{r}^{*}).$
2.2 Verifying Reasoning Steps
@@
The most prominent approach for assessing reasoning steps in both BoN and beam search test-time scaling to date has been PRMs (Lightman et al., 2024; Luo et al., 2024; Wang et al., 2024). PRMs are critics designed to estimate the quality of each partial reasoning state $\mathbf{r}_{1:t}$ by assigning a reward $R_{\text{PRM}}\!\left(\mathbf{r}_{1:t}\right)$ that reflects the likelihood of the chain eventually leading to a correct solution. PRMs are typically implemented by a separate LLM trained to evaluate the plausibility and correctness of intermediate reasoning steps. The annotation for training PRMs comes from various sources, including crowdsourcing, self-consistency checking, and synthetic data generation pipelines.
@@
In BS, the reward from a PRM is used as a step-wise quality function. In BoN, the score for the complete chain $\mathbf{r}^{(j)}$ can be obtained via a temporal aggregation of process rewards, e.g., minimum step score (Zhang et al., 2025c):
$Q_{\text{BoN}}\!\left(\mathbf{r}^{(j)}\right)\;=\;\min_{1\leq t\leq T^{(j)}}R_{\text{PRM}}\!\left(\mathbf{r}^{(j)}_{1:t}\right)$,
$Q_{\text{bs}}\!\left(\mathbf{r}^{(i,k)}_{1:t}\right)=R_{\text{PRM}}\left(\mathbf{r}^{(i,k)}_{1:t}\right)$.
@@
PRMs come with several limitations: (1) they are relatively large models, typically containing 1.5B-8B parameters, which leads to additional computational and memory overhead during inference; (2) PRMs are often domain-specific, being trained for tasks such as mathematical reasoning, and exhibit limited generalization to unseen domains.
Recent work also suggests using confidence scores for the verification of reasoning steps or entire reasoning trajectories as a drop-in replacement for PRMs (Fu et al., 2025): the lower the uncertainty, the more trustworthy is the step. This is motivated by their efficiency and generalization.
3 Verifying Reasoning Steps via Probing Internal States of LLMs
ReProbe model. We construct a reasoning probe as a plug-and-play module on top of a frozen LLM that outputs a probability of how likely the current reasoning step is correct. While PRMs rely solely on generated tokens, our probe introspectively leverages richer per-token features extracted from the base LLM’s internal states, which allows it to be both efficient and effective (see Fig 1).
The feature extractor $F(\mathbf{x},\mathbf{r}_{1:t-1},r_{t})$ can leverage various internal signals at each token position. Specifically, we investigate two sets of features: (1) Attn+Logit: attention weights over the 5 preceding tokens from all layers and logits of the top-K candidate generations; and (2) Hidden states from all layers. We select these feature sets because they have proven effective for factual hallucination detection (Shelmanov et al., 2025; Azaria and Mitchell, 2023). However, their role in reasoning supervision has been unexplored.
The architecture of ReProbe is as follows:
@@ 3 Verifying Reasoning Steps via Probing Internal States of LLMs
    Linear feature projection layer that adjusts the dimension of token-level features:
@@
    Transformer layers: a stack of $L$ layers that capture contextual dependencies:
@@
    Aggregation of token-level features $h_{L}^{\tau}$ into a step-level vector:
    $h_{r}=\frac{1}{|\mathbf{r}_{t}|}\sum_{\tau\in\mathbf{r}_{t}}h_{L}^{\tau}.$
@@
    Two-layer classification NN that outputs a final logit $\mathbf{z}$:
    $\mathbf{z}=\mathrm{Linear}_{2}\!\Big(\sigma\big(\mathrm{Dropout}(\mathrm{Linear}_{1}(h_{r}))\big)\Big),$
    where $\sigma(\cdot)$ is a GeLU activation function.
Reasoning step extraction is done in two ways. LLMs in non-thinking mode are instructed to generate each CoT step on a separate, self-contained line (see § F.1). LLMs reliably follow this format and produce well-formatted CoT steps, so they can be easily extracted during training and inference via simple pattern matching. LLMs operating in native thinking mode cannot be explicitly instructed to produce reasoning trajectories with clearly separated steps. To mitigate this problem, we treat each sentence as a reasoning step.
@@
Constructing training data. To construct the training data for ReProbe, we use 10.8K problems (prompts) from the training subset of PRM800K. This dataset is an established resource for training and evaluating LLM reasoning capabilities. As most PRMs are already trained on LLM generations derived from it, it enables a fair comparison (Wang et al., 2024; Zhang et al., 2025c).
We generate 3 reasoning trajectories per problem across 10.8K mathematical problems, yielding ~32K samples. We use nucleus sampling (top-k=50k{=}50, top-p=0.95p{=}0.95, temperature=1.0{=}1.0). For annotation and training efficiency, we restrict the length of the generation during training. This restriction is removed during evaluation to ensure that our results are free of length bias.
@@
We annotate the correctness of each reasoning step using an LLM as a judge. The judge is provided with the input question, the target LLM’s CoT steps and final answer, as well as the ground-truth answer, and is prompted to assess the correctness of each reasoning step. The annotation prompts are provided in  § F.2. We consider two approaches to step annotation: (1) an external verification setting, in which a larger LLM evaluates the reasoning steps – following Zheng et al. (2025), we use DeepSeek-R1; and (2) a self-supervised setting, in which the same LLM annotates its own generated CoT steps. Other details for reproducing the training data construction can be found in § B.1.
Training ReProbe. We train the probes using standard binary cross-entropy loss. Due to severe class imbalance, we adjust the class weights. Only the parameters of ReProbe are trained; the entire LLM body remains frozen. We further note that training ReProbes is highly cost-efficient: the total annotation cost is only \$200, and the required compute amounts to just 4 GH200 GPU-hours (see § B.3).
@@
Reasoning Probes (ReProbes)
$\bigstar$ ReProbe, Attn+Logit, Self-anno (Ours) 32K .529 .594 .260 .735 .779 .779 .394 .404 .461 .618 .559
$\bigstar$ ReProbe, Attn+Logit, DeepSeek-anno (Ours) 32K .465 .616 .243 .740 .802 .786 .395 .361 .441 .617 .551
$\bigstar$ ReProbe, Hidden States, Self-anno (Ours) 32K .558 .673 .264 .799 .819 .785 .381 .553 .498 .667 .604
$\bigstar$ ReProbe, Hidden States, DeepSeek-anno (Ours) 32K .534 .650 .281 .793 .817 .789 .344 .450 .488 .639 .582